# thanks: Corresponding author

### v1

@@ I Introduction
A significant number of researchers have dedicated efforts to investigating different attributes of these baryons, such as their masses and residues ShekariTousi:2024mso ; Ebert:2002ig ; Zhang:2008rt ; Wang:2010hs ; Lu:2017meb ; Rahmani:2020pol ; Yao:2018ifh ; Aliyev:2022rrf ; Aliev:2012iv ; Aliev:2019lvd ; Aliev:2012ru ; Padmanath:2019ybu ; Brown:2014ena ; Giannuzzi:2009gh ; Shah:2017liu ; Shah:2016vmd ; Yoshida:2015tia ; Li:2022ywz ; Wang:2010it ; Ortiz-Pacheco:2023kjn ; Wang:2018lhz ; Bagan:1992za ; Alrebdi:2022lat ; Wang:2010vn , chiral effective Lagrangian Qiu:2020omj , as well as the determination of strong coupling constants Olamaei:2021hjd ; Aliev:2021hqq ; Aliev:2020lly ; Alrebdi:2020rev ; Rostami:2020euc ; Olamaei:2020bvw ; Aliev:2020aon . Further studies have focused on their strong interaction dynamics and decay mechanisms Azizi:2020zin ; Qin:2021dqo ; Xiao:2017dly , radiative transitions Aliev:2021hqq ; Xiao:2017udy ; Lu:2017meb ; Rahmani:2020pol ; Li:2017pxa ; Ortiz-Pacheco:2023kjn , weak decay channels ShekariTousi:2025fjf ; Tousi:2024usi ; Gerasimov:2019jwp ; Wang:2017mqp ; Zhao:2018mrg ; Xing:2018lre ; Jiang:2018oak ; Gutsche:2019wgu ; Gutsche:2019iac ; Ke:2019lcf ; Cheng:2020wmk ; Hu:2020mxk ; Li:2020qrh ; Han:2021gkl ; Wang:2017azm ; Shi:2017dto ; Zhang:2018llc ; Ivanov:2020xmw ; Shi:2020qde ; Hu:2017dzi ; Li:2018epz ; Shi:2019hbf ; Shi:2019fph ; Sharma:2017txj ; Patel:2024mfn ; Gutsche:2017hux ; Gutsche:2018msz , magnetic properties Ozdem:2018uue ; Ozdem:2019zis , lifetimes Berezhnoy:2018bde , and mixing angles Aliev:2012nn , applying a variety of theoretical methods. To obtain precise predictions for these parameters, nonperturbative approaches such as the QCD sum rules formulated by Shifman, Vainshtein, and Zakharov in 1979 Shifman:1978bx ; Shifman:1978by play a crucial role. This framework employs the QCD Lagrangian and relies on correlation functions constructed with different interpolating currents. The QCD sum rule method has yielded numerous reliable estimates for hadronic features, corroborated by experimental findings, and remains a powerful tool in hadron physics Aliev:2010uy ; Aliev:2009jt ; Aliev:2012ru ; Agaev:2016dev ; Azizi:2016dhy .
@@
On the other hand, the study of the mass of the doubly heavy spin-$\frac{3}{2}$ baryons is done in the literature. The ground state mass determination for these baryons have been carried out using a variety of theoretical frameworks, including lattice QCD simulations Brown:2014ena , constituent quark models Yoshida:2015tia , relativistic quark models Ebert:2002ig , the QCD sum rules approach Aliev:2012iv ; Bagan:1992za ; Zhang:2008rt ; Wang:2010vn , hypercentral constituent quark models Shah:2016vmd ; Shah:2017liu , and solutions to the Salpeter equation Giannuzzi:2009gh . To improve the precision and minimize uncertainties, we expand the analysis to include nonperturbative operators with dimensions reaching ten. Earlier studies using sum rules Aliev:2012iv were limited to nonperturbative operators up to dimension five.
@@
Various theoretical methods have been employed to analyze the mass of doubly heavy baryons with spin-$\frac{3}{2}$ in the $1P$ excited configuration, including the hypercentral constituent quark model Shah:2016vmd ; Shah:2017liu , constituent quark models Yoshida:2015tia , and the QCD sum rules Alrebdi:2022lat ; Wang:2010it . The mass of $2S$ excited state of these baryons has also been computed in various frameworks such as the hypercentral constituent quark model Shah:2016vmd ; Shah:2017liu , constituent quark models Yoshida:2015tia , and the QCD sum rule technique Alrebdi:2022lat . In this work, we aim to achieve improved precision in calculating the masses of all baryon members in these excited states by incorporating higher dimensional nonperturbative contributions, as outlined above. We calculate the residues of doubly heavy spin-$\frac{3}{2}$ baryons across the ground state, $1P$ and $2S$ excited states, with a high accuracy and up to ten dimensional nonperturbative operator in the operator product expansion (OPE) as well. Previous research has limited accuracy and it is up to five dimensions in OPE: The ground states have been analyzed using the QCD sum rule method in Refs. Aliev:2012iv ; Wang:2010vn ; Bagan:1992za . The residues corresponding to the $1P$ and $2S$ states were determined using QCD sum rules, as detailed in Refs. Alrebdi:2022lat ; Wang:2010it and Ref. Alrebdi:2022lat , respectively. We extend these studies by focusing on the mass spectra and residues for the spin-$\frac{3}{2}$ doubly heavy baryons in the ground state as well as the initial radial and orbital excited configurations, which serve as essential inputs for investigating the decay mechanisms and interaction dynamics involving such baryons. The findings may serve as valuable input for upcoming experimental investigations of such baryons at particle physics facilities like the LHC. The prospect of soon detecting new doubly heavy baryons motivates theoretical explorations that could significantly advance our knowledge of heavy quark dynamics within hadrons. These investigations serve as a powerful means to bridge the gap between perturbative and nonperturbative QCD, providing essential theoretical groundwork for future experimental findings. Furthermore, doubly heavy baryons research may unlock deeper insights into the fundamental nature of other heavy baryon systems.
@@ II.2 QCD side
where $S_{q}$ and $S_{Q}$ are defined as the light and heavy quark propagators, respectively, and we also define $\widetilde{S}$ as $CS^{T}C$. The propagators of light and heavy quarks in coordinate space are as follows:
@@
Subsequently, with the mass and residue of the ground state determined, the threshold parameter is increased to introduce the first orbital excitation, $1P$. In this instance, the initial two terms on the left-hand side of Eq. (15), corresponding to the $1S$ and $1P$ configurations, are taken into account. Consequently, the third term, representing the radially excited, $2S$ state, is incorporated within the continuum, thereby implementing the ground state + first orbitally excited state + continuum scheme. Given the established mass and residue values for the ground state, two novel unknowns emerge, namely the mass and residue of the $1P$ state. These can be ascertained analogously to the ground state determination. Ultimately, the ground state + first orbitally excited state + first radially excited state + continuum scheme is employed to deduce the mass and residue of the radially excited $2S$ state, achieved by further incrementing the continuum threshold and utilizing the spectroscopic parameters of the ground and $1P$ states as inputs. These methodologies will be revisited in the subsequent section, considering pertinent numerical analysis.
@@ III Numerical results and discussions
Now, we detail the numerical evaluation of our findings concerning the masses and residues of doubly heavy spin-$(\frac{3}{2})$} baryons, encompassing their ground, first orbitally, and first radially excited states. We incorporate several input parameters, as delineated in table 2. In these calculations, the masses of the $u$ and $d$ quarks are neglected, whereas the mass of the $s$ quark is considered. As evidenced by table 2, we present condensate values up to mass dimension six, though the calculations extend to mass dimension ten. The nonperturbative operators exceeding mass dimension six (specifically, dimensions seven through ten) are expressed in terms of lower-dimensional operators, leveraging the QCD factorization hypothesis. This procedure is implemented during the computation of heavy and light quark propagators, as presented in the preceding section. Upon multiplying the propagators via Eq. (II.2), the higher-dimensional operators manifest as products of operators with lower dimensions, with their values derived from the numerical data presented in table 2.
Parameters Values
$m_{c}$ $1.27\pm 0.02~\mathrm{GeV}$ Zyla:2020zbs
$m_{b}$ $4.18^{+0.03}_{-0.02}~\mathrm{GeV}$ Zyla:2020zbs
$m_{s}$ $93^{+11}_{-5}~\mathrm{MeV}$ Zyla:2020zbs
$\langle\bar{q}q\rangle$ $(-0.24\pm 0.01)^{3}$ $\mathrm{GeV}^{3}$ Belyaev:1982sa
$\langle\bar{s}s\rangle$ $0.8\langle\bar{q}q\rangle$ Belyaev:1982sa
$m_{0}^{2}$ $(0.8\pm 0.1)$ $\mathrm{GeV}^{2}$ Belyaev:1982sa
$\langle\overline{q}g_{s}\sigma Gq\rangle$ $m_{0}^{2}\langle\bar{q}q\rangle$ Belyaev:1982sa
$\langle g_{s}^{2}G^{2}\rangle$ $4\pi^{2}(0.012\pm 0.004)$ $\mathrm{GeV}^{4}$Belyaev:1982cd
@@
$\langle\frac{\alpha_{s}}{\pi}G^{2}\rangle$ $(0.012\pm 0.004)$ $\mathrm{GeV}^{4}$Belyaev:1982cd
$\langle g_{s}^{3}G^{3}\rangle$ $(0.57\pm 0.29)$ $\mathrm{GeV}^{6}$Narison:2015nxh
Table 2: Key parameters utilized in the computations.
@@
Apart from the parameters listed in Table 2, the QCD sum rules require two additional parameters that need to be fixed: the Borel parameter, $M^{2}$, and the threshold parameter, $s_{0}$. These results are obtained by applying the standard conditions of the QCD sum rule approach, which include minimal dependence on auxiliary parameters, dominance of the pole contribution, and the convergence of the OPE. OPE convergence is attained when the perturbative contribution predominates over the nonperturbative components, with higher-dimensional nonperturbative operators exhibiting progressively diminishing contributions. Pole dominance ensures that the resonances under investigation obtain a primary contribution relative to higher excited states and the continuum. Consequently, the upper bound of the $M^{2}$ parameter is dictated by the stipulation that the pole contributions (PC) exceed those originating from higher excited states and the continuum. According to these considerations, the subsequent conditions must be satisfied:
@@
The continuum threshold parameter $s_{0}$, which reflects the onset of excited state contributions, is determined to ensure that the ground state dominates within the chosen analysis window. This selection prevents contamination from higher resonances and continuum states in the sum rule evaluations. The chosen intervals for both the Borel mass parameter $M^{2}$ and the threshold value $s_{0}$ for each channel are listed in table 3. The numerical analysis reveals that the physical quantities remain largely stable under changes in the auxiliary parameters within the chosen intervals for the borel mass parameter $M^{2}$ and continuum threshold $s_{0}$. To illustrate the dependence of the masses on these parameters, we provide Figs. 2, 3, and 4. The findings indicate robust stability concerning the Borel parameter and continuum threshold across their operational ranges. Among the analyzed states, the $\Xi_{cc}^{*}$ baryon shows representative behavior, and thus, its graphical results are included as a demonstration. The main sources of uncertainty in the computed values arise from the uncertainties in auxiliary and input parameters.

### v2

@@ I Introduction
A significant number of researchers have dedicated efforts to investigating different attributes of these baryons, such as their masses and residues ShekariTousi:2024mso ; Ebert:2002ig ; Zhang:2008rt ; Wang:2010hs ; Lu:2017meb ; Rahmani:2020pol ; Yao:2018ifh ; Aliyev:2022rrf ; Aliev:2012iv ; Aliev:2019lvd ; Aliev:2012ru ; Padmanath:2019ybu ; Brown:2014ena ; Giannuzzi:2009gh ; Shah:2017liu ; Shah:2016vmd ; Yoshida:2015tia ; Li:2022ywz ; Wang:2010it ; Ortiz-Pacheco:2023kjn ; Wang:2018lhz ; Bagan:1992za ; Alrebdi:2022lat ; Wang:2010vn ; ShekariTousi:2026efp , chiral effective Lagrangian Qiu:2020omj , as well as the determination of strong coupling constants Olamaei:2021hjd ; Aliev:2021hqq ; Aliev:2020lly ; Alrebdi:2020rev ; Rostami:2020euc ; Olamaei:2020bvw ; Aliev:2020aon . Further studies have focused on their strong interaction dynamics and decay mechanisms Azizi:2020zin ; Qin:2021dqo ; Xiao:2017dly , radiative transitions Aliev:2021hqq ; Xiao:2017udy ; Lu:2017meb ; Rahmani:2020pol ; Li:2017pxa ; Ortiz-Pacheco:2023kjn , weak decay channels ShekariTousi:2025fjf ; Tousi:2024usi ; Gerasimov:2019jwp ; Wang:2017mqp ; Zhao:2018mrg ; Xing:2018lre ; Jiang:2018oak ; Gutsche:2019wgu ; Gutsche:2019iac ; Ke:2019lcf ; Cheng:2020wmk ; Hu:2020mxk ; Li:2020qrh ; Han:2021gkl ; Wang:2017azm ; Shi:2017dto ; Zhang:2018llc ; Ivanov:2020xmw ; Shi:2020qde ; Hu:2017dzi ; Li:2018epz ; Shi:2019hbf ; Shi:2019fph ; Sharma:2017txj ; Patel:2024mfn ; Gutsche:2017hux ; Gutsche:2018msz , magnetic properties Ozdem:2018uue ; Ozdem:2019zis , lifetimes Berezhnoy:2018bde , and mixing angles Aliev:2012nn , applying a variety of theoretical methods. To obtain precise predictions for these parameters, nonperturbative approaches such as the QCD sum rules formulated by Shifman, Vainshtein, and Zakharov in 1979 Shifman:1978bx ; Shifman:1978by play a crucial role. This framework employs the QCD Lagrangian and relies on correlation functions constructed with different interpolating currents. The QCD sum rule method has yielded numerous reliable estimates for hadronic features, corroborated by experimental findings, and remains a powerful tool in hadron physics Aliev:2010uy ; Aliev:2009jt ; Aliev:2012ru ; Agaev:2016dev ; Azizi:2016dhy .
@@
On the other hand, the study of the mass of the doubly heavy spin-$\frac{3}{2}$ baryons is done in the literature. The ground state mass determination for these baryons have been carried out using a variety of theoretical frameworks, including lattice QCD simulations Brown:2014ena , constituent quark models Yoshida:2015tia , relativistic quark models Ebert:2002ig , the QCD sum rules approach Aliev:2012iv ; Bagan:1992za ; Zhang:2008rt ; Wang:2010vn , hypercentral constituent quark models Shah:2016vmd ; Shah:2017liu , and solutions to the Salpeter equation Giannuzzi:2009gh . To improve the precision and minimize uncertainties, we expand the analysis to include nonperturbative operators with dimensions reaching ten. Earlier studies using sum rules Aliev:2012iv ; Bagan:1992za ; Zhang:2008rt ; Wang:2010vn were limited to nonperturbative operators up to dimension five and six.
@@
Various theoretical methods have been employed to analyze the mass of doubly heavy baryons with spin-$\frac{3}{2}$ in the $1P$ excited configuration, including the hypercentral constituent quark model Shah:2016vmd ; Shah:2017liu , constituent quark models Yoshida:2015tia , and the QCD sum rules Alrebdi:2022lat ; Wang:2010it . The mass of $2S$ excited state of these baryons has also been computed in various frameworks such as the hypercentral constituent quark model Shah:2016vmd ; Shah:2017liu , constituent quark models Yoshida:2015tia , and the QCD sum rule technique Alrebdi:2022lat . In this work, we aim to achieve improved precision in calculating the masses of all baryon members in these excited states by incorporating higher-dimensional nonperturbative contributions, as outlined above. We calculate the residues of doubly heavy spin-$\frac{3}{2}$ baryons across the ground state, $1P$ and $2S$ excited states, with a high accuracy and up to ten dimensional nonperturbative operator in the operator product expansion (OPE) as well. The ground states have been analyzed using the QCD sum rule method in Refs. Aliev:2012iv ; Wang:2010vn ; Bagan:1992za . The residues corresponding to the $1P$ and $2S$ states were determined using QCD sum rules, as detailed in Refs. Alrebdi:2022lat ; Wang:2010it and Ref. Alrebdi:2022lat , respectively. We extend these studies by focusing on the mass spectra and residues for the spin-$\frac{3}{2}$ doubly heavy baryons in the ground state as well as the initial radial and orbital excited configurations, which serve as essential inputs for investigating the decay mechanisms and interaction dynamics involving such baryons. The findings may serve as valuable input for upcoming experimental investigations of such baryons at particle physics facilities like the LHC. The prospect of soon detecting new doubly heavy baryons motivates theoretical explorations that could significantly advance our knowledge of heavy quark dynamics within hadrons. These investigations serve as a powerful means to bridge the gap between perturbative and nonperturbative QCD, providing essential theoretical groundwork for future experimental findings. Furthermore, doubly heavy baryons research may unlock deeper insights into the fundamental nature of other heavy baryon systems.
@@ II.2 QCD side
where $S_{q}$ and $S_{Q}$ are defined as the light and heavy quark propagators, respectively, and we also define $\widetilde{S}$ as $CS^{T}C$. The propagators of light and heavy quarks in coordinate space are as follows Reinders:1984sr :
@@
Subsequently, once the mass and residue of the ground state have been determined with sufficient accuracy, the analysis proceeds by gradually increasing the continuum threshold $s_{0}$ so as to accommodate the contribution of the first orbital excitation, $1P$. In this instance, the initial two terms on the left-hand side of Eq. (15), corresponding to the $1S$ and $1P$ configurations, are taken into account. Consequently, the third term, representing the radially excited, $2S$ state, is incorporated within the continuum, thereby implementing the ground state + first orbitally excited state + continuum scheme. Within this approach, the mass and residue of the ground state serve as known inputs, while two additional unknowns appear specifically, the mass and residue of the $1P$ state. These new parameters can then be extracted in close analogy to the procedure already employed for the ground state. Ultimately, the ground state + first orbitally excited state + first radially excited state + continuum scheme is employed to deduce the mass and residue of the radially excited $2S$ state, achieved by further incrementing the continuum threshold $s_{0}$ and utilizing the spectroscopic parameters of the ground and $1P$ states as inputs. These stepwise methodologies progressively extending the spectrum from the ground state to the first orbital excitation and subsequently to the first radial excitation will be systematically revisited in the next section, where detailed numerical analyses are provided to substantiate the outlined procedure.
@@ III Numerical results and discussions
Now, we detail the numerical evaluation of our findings concerning the masses and residues of doubly heavy spin-$(\frac{3}{2})$} baryons, encompassing their ground, first orbitally, and first radially excited states.
In this work, the quark masses are adopted within the $\overline{MS}$ scheme using the values reported by the Particle Data Group (PDG) Zyla:2020zbs : $m_{c}(m_{c})=(1.27\pm 0.02)\,\mathrm{GeV}$, $m_{b}(m_{b})=(4.18\pm 0.03)\,\mathrm{GeV}$, and $m_{s}(\mu=2\,\mathrm{GeV})=(0.095\pm 0.005)\,\mathrm{GeV}$. The variation of these masses with respect to the energy scale is determined through the renormalization group equations. In our analysis, the vacuum condensates are taken at the typical reference scale $\mu=1\,\mathrm{GeV}$ with the standard values $\langle\bar{q}q\rangle=-(0.24\pm 0.01\,\mathrm{GeV})^{3}$, $\langle\bar{s}s\rangle=0.8\langle\bar{q}q\rangle$, $\langle\bar{q}g_{s}\sigma Gq\rangle=m_{0}^{2}\langle\bar{q}q\rangle$, $m_{0}^{2}=(0.8\pm 0.1)\,\mathrm{GeV}^{2}$, $\langle\frac{\alpha_{s}GG}{\pi}\rangle=(0.012\pm 0.004)\,\mathrm{GeV}^{4}$ and $\langle g_{s}^{3}G^{3}\rangle=(0.57\pm 0.29)$ $\mathrm{GeV}^{6}$, as commonly adopted in QCD sum rule calculations Belyaev:1982sa ; Belyaev:1982cd ; Narison:2015nxh . To account for the scale dependence of the relevant parameters, we apply the renormalization group evolution laws (see for instance Wang:2021dcb ; Wang:2014oca and references therein):
$m_{s}(\mu)$ $=$ $m_{s}(2\,\mathrm{GeV})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(2\,\mathrm{GeV})}\right]^{\frac{4}{9}}\,$,
$m_{c}(\mu)$ $=$ $m_{c}(m_{c})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(m_{c})}\right]^{\frac{12}{25}}\,$,
$m_{b}(\mu)$ $=$ $m_{b}(m_{b})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(m_{b})}\right]^{\frac{12}{23}}\,$,
$\langle\bar{q}q\rangle(\mu)$ $=$ $\langle\bar{q}q\rangle(1\,\mathrm{GeV})\left[\frac{\alpha_{s}(1\,\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{12}{33-2n_{f}}}\,$,
@@
$\langle\bar{q}g_{s}\sigma Gq\rangle(\mu)$ $=$ $\langle\bar{q}g_{s}\sigma Gq\rangle(1\,\mathrm{GeV})\left[\frac{\alpha_{s}(1\,\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{2}{33-2n_{f}}}\,$,
$\alpha_{s}(\mu)$ $=$ $\frac{1}{b_{0}t}\left[1-\frac{b_{1}}{b_{0}^{2}}\frac{\ln t}{t}+\frac{b_{1}^{2}(\ln^{2}t-\ln t-1)+b_{0}b_{2}}{b_{0}^{4}t^{2}}\right]\,$, (19)
where $t=\ln\left(\frac{\mu^{2}}{\Lambda^{2}}\right)$, and the coefficients take the values $b_{0}=\frac{33-2n_{f}}{12\pi}$, $b_{1}=\frac{153-19n_{f}}{24\pi^{2}}$, and $b_{2}=\frac{2857-\frac{5033}{9}n_{f}+\frac{325}{27}n_{f}^{2}}{128\pi^{3}}$. The QCD scale parameter $\Lambda$ is taken to be $213\,\mathrm{MeV}$, $296\,\mathrm{MeV}$, and $339\,\mathrm{MeV}$ for the flavor numbers $n_{f}=5$, $4$, and $3$, respectively, following PDG conventions  Zyla:2020zbs . In these calculations, the masses of the $u$ and $d$ quarks are neglected. All input parameters are therefore evolved to an appropriate energy scale $\mu$ consistent with this choice in order to determine the corresponding hadron masses and residues. Here, we take $n_{f}=4$ for the $\Xi_{cc}^{*}$ and $\Omega_{cc}^{*}$ baryons and $n_{f}=5$ for the $\Xi_{bc}^{*}$, $\Xi_{bb}^{*}$, $\Omega_{bc}^{*}$ and $\Omega_{bb}^{*}$ baryons. As we have the bottom and charm quarks in the considered states, we take the interval $\mu$ =(2-4) GeV for these calculations. Finally, we report the averaged values of the mass and the residue obtained for $\mu$= 2, 3 and 4 GeV. The gluon condensate values $\langle\frac{\alpha_{s}GG}{\pi}\rangle=(0.012\pm 0.004)\,\mathrm{GeV}^{4}$ and $\langle g_{s}^{3}G^{3}\rangle=(0.57\pm 0.29)$ $\mathrm{GeV}^{6}$ remain widely used in contemporary QCD sum rule studies, as it continues to represent the most reliable estimate in the absence of more precise determinations.
In the present work, the operator product expansion (OPE) is carried out consistently up to vacuum condensates of dimension 10. The higher-dimensional condensates are expressed in terms of lower-dimensional ones using the factorization assumption. As discussed by Shifman, Vainshtein, Zakharov, and Ioffe Shifman:1978bx ; Shifman:1978by ; Ioffe:2005ym , the factorization (or vacuum saturation) hypothesis is expected to hold with an accuracy of order $1/N_{c}^{2}$. In the theoretical limit of large $N_{c}$, the correction term $1/N_{c}^{2}$ tends to zero, indicating that the approximation becomes exact. Nevertheless, for the physical case of $N_{c}=3$, the suppression factor is finite, $1/N_{c}^{2}\approx 10\%$, which defines the typical uncertainty associated with factorization. Consequently, violations of factorization represent an intrinsic source of theoretical uncertainty in QCD sum rule calculations involving multi-quark condensates. To parametrize possible non-factorizable contributions in a systematic and transparent way, we introduce a phenomenological parameter $\kappa$ that measures deviations from vacuum saturation  Wang:2021dcb . In particular, the factorized condensate structures are modified as
$\langle\bar{q}q\rangle^{2}$ $\rightarrow\kappa\,\langle\bar{q}q\rangle^{2},$
$\langle\bar{q}q\rangle\left\langle\frac{\alpha_{s}}{\pi}GG\right\rangle$ $\rightarrow\kappa\,\langle\bar{q}q\rangle\left\langle\frac{\alpha_{s}}{\pi}GG\right\rangle,$
$\langle\bar{q}q\rangle\langle\bar{q}g_{s}\sigma Gq\rangle$ $\rightarrow\kappa\,\langle\bar{q}q\rangle\langle\bar{q}g_{s}\sigma Gq\rangle,$
$\left\langle\frac{\alpha_{s}}{\pi}GG\right\rangle\langle\bar{q}g_{s}\sigma Gq\rangle$ $\rightarrow\kappa\,\left\langle\frac{\alpha_{s}}{\pi}GG\right\rangle\langle\bar{q}g_{s}\sigma Gq\rangle,$
$\langle\bar{q}g_{s}\sigma Gq\rangle^{2}$ $\rightarrow\kappa\,\langle\bar{q}g_{s}\sigma Gq\rangle^{2},$
$\langle\bar{q}q\rangle^{2}\left\langle\frac{\alpha_{s}}{\pi}GG\right\rangle$ $\rightarrow\kappa\,\langle\bar{q}q\rangle^{2}\left\langle\frac{\alpha_{s}}{\pi}GG\right\rangle,\cdots.$ (20)
Here, $\kappa=1$ corresponds to exact vacuum saturation, while $\kappa>1$ accounts for possible non-factorizable effects. Several determinations of the parameter $\kappa$ have been reported in the literature, with its magnitude varying according to the specific channel under investigation. The available estimates span a broad interval, starting from $\kappa=1$ Ioffe:2010zz , increasing to approximately 2233 Narison:2002woh , and reaching values as large as $6$ in certain analyses Leinweber:1995fn . Although, in principle, different condensate structures may be associated with different $\kappa$ values, introducing independent parameters would significantly enlarge the parameter space without clear phenomenological guidance. For this reason, and to maintain a controlled uncertainty analysis, we adopt a common $\kappa$ for all factorized higher-dimensional condensates. To obtain a conservative and realistic estimate of the theoretical uncertainty induced by the factorization assumption, we vary the parameter within the range
$1\leq\kappa\leq 5.$
For each value of $\kappa$, the sum rules are recalculated, and the corresponding masses and residues are extracted. The maximal deviation of these quantities relative to their central values is then incorporated into the total theoretical uncertainty. The calculations resulted in an estimated uncertainty of roughly $2\%$ in the determined mass and residue of the hadrons.
@@
Apart from these parameters, the QCD sum rules require two additional parameters that need to be fixed: the Borel parameter, $M^{2}$, and the threshold parameter, $s_{0}$. These results are obtained by applying the standard conditions of the QCD sum rule approach, which include minimal dependence on auxiliary parameters, dominance of the pole contribution, and the convergence of the OPE. OPE convergence is attained when the perturbative contribution predominates over the nonperturbative components, with higher-dimensional nonperturbative operators exhibiting progressively diminishing contributions. Pole dominance ensures that the resonances under investigation obtain a primary contribution relative to higher excited states and the continuum. Consequently, the upper bound of the $M^{2}$ parameter is dictated by the stipulation that the pole contributions (PC) exceed those originating from higher excited states and the continuum. According to these considerations, the subsequent conditions must be satisfied:
@@
The continuum threshold parameter $s_{0}$, which reflects the onset of excited state contributions, is determined to ensure that the ground state dominates within the chosen analysis window. This selection prevents contamination from higher resonances and continuum states in the sum rule evaluations. The chosen intervals for both the Borel mass parameter $M^{2}$ and the threshold value $s_{0}$ for each channel are listed in table 2. The numerical analysis reveals that the physical quantities remain largely stable under changes in the auxiliary parameters within the chosen intervals for the borel mass parameter $M^{2}$ and continuum threshold $s_{0}$. To illustrate the dependence of the masses on these parameters, we provide Figs. 2, 3, and 4 for $\kappa=1$ and at average value of the scale $\mu$. The findings indicate robust stability concerning the Borel parameter and continuum threshold across their operational ranges. Among the analyzed states, the $\Xi_{cc}^{*}$ baryon shows representative behavior, and thus, its graphical results are included as a demonstration. The main sources of uncertainty in the computed values arise from the uncertainties in auxiliary and input parameters.